# Pref-GUIDE: Continual Policy Learning from Real-Time Human Feedback via Preference-Based Learning

### v1

@@ 4.1 Inspiration: Human Feedback is Inconsistent Over Time
To validate our hypothesis, we processed each human evaluator’s $\mathcal{D}_{\text{real-time}}$ by encoding the trajectories using a pretrained image encoder, DINO Oquab et al. (2023), followed by a t-SNE Van der Maaten & Hinton (2008) projection for visualization. The colors represent different values of the human feedback. As shown in Figure 3, there is no clear alignment between similar trajectory embeddings and scalar feedback. This result supports our hypothesis that feedback criteria shift as humans adjust their expectations during guidance. Visualizations from other evaluators are provided in the Appendix A.1.
@@ 6 Conclusion
We introduced Pref-GUIDE, a method that converts real-time scalar human feedback into preference-based data to train reward models for continual policy learning. Pref-GUIDE Individual mitigates temporal inconsistency by transforming continuous feedback labels into moving-window sampled preference labels. Pref-GUIDE Voting enhances robustness by aggregating reward models across individuals into a consensus-driven signal. Across three visual RL environments, Pref-GUIDE Individual outperforms regression-based baselines such as GUIDE when feedback is of high quality, and Pref-GUIDE Voting maintains strong performance even under noisy conditions, surpassing both GUIDE and expert-designed dense rewards. Together, these results highlight the scalability and stability of structured preference learning from real-time human input. Some future directions can take Pref-GUIDE to study the incorporation of richer multimodal human feedback and adapting it to broader domains such as robotics and multi-agent settings.
7 Acknowledgment
This work is supported by the ARL STRONG program under awards W911NF2320182, W911NF2220113, and W911NF2420215, and by gift supports from BMW and OpenAI.

### v2

@@ 4.1 Inspiration: Human Feedback is Inconsistent Over Time
To validate our hypothesis, we processed each human evaluator’s $\mathcal{D}_{\text{real-time}}$ by encoding the trajectories using a pretrained image encoder, DINO Oquab et al. (2023), followed by a t-SNE Van der Maaten & Hinton (2008) projection for visualization. The colors represent different values of the human feedback. As shown in Figure 3, there is no clear alignment between similar trajectory embeddings and scalar feedback. This trend is less obvious in Bowling due to its relatively shorter horizon and rather static observations, compared with Find Treasure and Hide and Seek 1v1. However, it’s clear in the other tasks that require longer horizons and explorations, human feedback is not consistent over time. This result supports our hypothesis that feedback criteria shift as humans adjust their expectations during guidance. Visualizations from other evaluators are provided in the Appendix A.1.
@@ 6 Conclusion
We introduced Pref-GUIDE, a method that converts real-time scalar human feedback into preference-based data to train reward models for continual policy learning. Pref-GUIDE Individual mitigates temporal inconsistency by transforming continuous feedback labels into moving-window sampled preference labels. Pref-GUIDE Voting enhances robustness by aggregating reward models across individuals into a consensus-driven signal. Across three visual RL environments, Pref-GUIDE Individual outperforms regression-based baselines such as GUIDE when feedback is of high quality, and Pref-GUIDE Voting maintains strong performance even under noisy conditions, surpassing both GUIDE and expert-designed dense rewards. Together, these results highlight the scalability and stability of structured preference learning from real-time human input.
As for the limitation of our method, for the two design choices of Pref-GUIDE Individual: no-preference range and moving window, we assumed that the same set of parameters is sufficient to model different human evaluators. Though our results showed that the same set of parameters performed well, these parameter selections may be further improved. For instance, more optimal subject-specific no-preference ranges and moving window sizes could potentially be inferred through a warm-up phase of human guidance. Furthermore, for every human evaluator, we trained a separate model. This model is rather small, and the additional training overhead is minimal. However, it is possible that such a model can add more training overhead when the reward model size needs to be large. Exploring more efficient reward models can be an interesting future direction. For Pref-GUIDE Voting, more algorithm improvements or more sophisticated aggregation methods can also be studied to incorporate the human guidance from all subjects.
@@ A.1 Visualization of Trajectories vs Human Feedback
Furthermore, we also visualize the trajectories that correspond to the neighbor points in the embedding space to further verify our hypothesis. In Figure 10,  11, and  12, for each task, we uniformly selected 8 anchor trajectories, and found the closest trajectory to every one of them in t-SNE space that is at least 100 timesteps apart to avoid temporal similarity. We visualized these trajectories along with the corresponding human feedback. As shown in the resulting trajectories, despite that they are semantically and visually very similar or near identical within the same trajectory, the human feedback is very distinct at different timesteps. These observations show strong evidence that human feedback is not consistent.
Figure 10: Visualization of Bowling from Subject 0. The trajectories that are semantically or visually similar to each other at different timesteps receive very distinct human feedback values.
Figure 11: Visualization of Find Treasure from Subject 0. The trajectories that are semantically or visually similar to each other at different timesteps receive very distinct human feedback values.
Figure 12: Visualization of Hide and Seek 1v1 from Subject 0. The trajectories that are semantically or visually similar to each other at different timesteps receive very distinct human feedback values.
@@ A.3 Pref-GUIDEVoting Algorithm
A.4 Experiment Details
Each experimental result was averaged over 5 seeds. The dense rewards for Hide and Seek 1v1 and Find Treasure were defined based on the visibility of the treasure or hider and the distance between the agent and the target.
A.5 Hardware Requirement
We conducted the experiments using NVIDIA A100 and NVIDIA RTX A6000. However, our experiment can be run on a single desktop with a single GPU, such as NVIDIA RTX 4070.